# Designing a skilled soccer team for RoboCup: exploring Skill-Set-Primitives through reinforcement learning

### v1

@@ Abstract
The RoboCup 3D Soccer Simulation League serves as a competitive platform for showcasing innovation in autonomous humanoid robot agents through simulated soccer matches. Our team, FC Portugal, developed a new codebase from scratch in Python after RoboCup 2021. The team’s performance is based on a set of skills centered around novel unifying primitives and a custom, symmetry-extended version of the Proximal Policy Optimization algorithm. Our methods have been thoroughly tested in official RoboCup matches, where FC Portugal has won the last two main competitions, in 2022 and 2023. This paper presents our training framework, as well as a timeline of skills developed using our skill-set-primitives, which considerably improve the sample efficiency and stability of skills, and motivate seamless transitions. We start with a significantly fast sprint-kick developed in 2021 and progress to the most recent skill set, which includes a multi-purpose omnidirectional walk, a dribble with unprecedented ball control, a solid kick, and a push skill. The push tackles both low-level collision-prone scenarios and high-level strategies to increase ball possession. We address the resource-intensive nature of this task through an innovative multi-agent learning approach. Finally, we release the codebase of our team to the RoboCup community, enabling other teams to transition to Python more easily and providing new teams with a robust and modern foundation upon which they can build new features.
@@ 1 Introduction
RoboCup is an international competition standing at the forefront of robotics research and innovation. The subject of our work is the 3D Soccer Simulation League (3DSSL) — a simulated soccer competition involving two teams of 11 autonomous humanoid robots. The rapid advancement of reinforcement learning (RL) techniques has paved the way for the creation of highly skilled robotic players, capable of efficiently controlling the ball while demonstrating tactical understanding to cooperate with teammates.
One of the main challenges lies in creating a cohesive structure of skills that works in harmony at different levels of abstraction — from motor control to strategy. This paper presents a novel approach to develop that structure using primitives that represent the commonalities of a skill set. All the methods detailed in this paper have been validated through participation in several official competitions, serving as the central pillar of our team, FC Portugal.
Figure 1: Timeline of skills (based on skill-set-primitives) that were used in official competitions. Each rhombus denotes the tally of goals directly attributed to each skill in the respective competition
Figure 1 depicts a timeline of skills developed with skill-set-primitives. In 2020, we developed the Sprint-Kick  [^1] First Sprint-Kick demonstration in 2020 https://youtu.be/Yy1yCM5hwZI, a behavior that sprinted toward the ball during kickoff, resulting in 22% of our team’s goals in the Portuguese Robotics Open (PRO) 2021. In RoboCup (RC) 2021 (virtual event), the kickoff was refined, accounting for 38% of all our goals  [^2] Successful kickoffs in RoboCup 2021 https://youtu.be/3MND8RVUPBQ. We secured 3rd and 4th place, respectively, having good individual skills but lacking cohesion.
@@
The dominant dribbling skill allowed any player, even the goalkeeper, to quickly bypass all opponents and score, while keeping the ball between both feet, making defense nearly impossible without fouling. To improve gameplay, RoboCup 2023 (Bordeaux) introduced new rules to prevent ball-holding  [^3] The official rules can be found at https://ssim.robocup.org/3d-simulation/3d-rules/, rendering our Dribble and Push skills unusable. Due to time constraints, only the Dribble was retrained to comply with the new rules. FC Portugal won the main competition for the third time in 2023.
@@ 2.2 Learning algorithm
The reinforcement learning algorithm used in this work is the Proximal Policy Optimization (PPO) [44] extended with Proximal Symmetry Loss (PSL) [^4] PPO+PSL implementation: https://github.com/m-abr/Adaptive-Symmetry-Learning [5]. PPO was chosen due to its good performance in single-agent and cooperative multi-agent games [59]. PSL leverages the symmetry properties of the NAO robot (sagittal plane), to improve the training sample efficiency and also motivate a more human-like behavior, without impeding asymmetric exploration. The extended objective function can be written as:
@@
[…]
@@
[…]
@@
The RoboCup 3D simulation league uses SimSpark [58] as its physical multi-agent simulator. The league’s environment is a 30 m by 20 m soccer field containing several landmarks that can be used for self-localization: goalposts, corner flags, and lines. Each team consists of 11 humanoid robots modeled after the NAO robot, created by SoftBank Robotics. Agents get internal data (joints, accelerometer, gyroscope, and foot pressure sensors) with a 1-step delay every 0.02 s, and visual data (restricted to a ${}^{\circ}$ vision cone) every 0.04 s (the visual data interval was 0.06 s prior to 2023). Agents can send joint commands every 0.02 s and message teammates every 0.04 s. In addition to the standard NAO model (type 0), there are alternative types with longer limbs (type 1), quicker feet (type 2), wider hips and longest limbs (type 3), and added toes (type 4).
@@ 3 Related Work
RoboCup is an international robotics competition that drives AI and robotics research, featuring standardized challenges for robotic teams. Most of the original challenges in 1997 [27] still stand today in the 3D soccer simulation league (3DSSL), including multiagent collaboration, real-time reasoning, handling dynamic environments, and learning complex tasks.
@@ 3.1 Learning paradigms
Developing a team involves the creation of an interplay of skills at different levels of abstraction, requiring seamless transitions and coordination between objectives of varying granularity. Hierarchical reinforcement learning is one way of tackling this issue by allowing the decomposition of complex tasks. In light of this, Stone et al. [53] proposed the Layered Learning (LL) paradigm, where tasks are learned in a bottom-up approach, originally in a sequential way. A lower-level layer (A) is learned and frozen before learning the next layer (B). Posterior formulations include concurrent [57] and overlapping variations [33; 34]. In concurrent LL, (A) is initially learned in isolation, being later refined while simultaneously learning (B). In overlapping LL, there are 3 paradigms that can be succinctly described as: the combination of independently learned layers, a partial version of concurrent LL where (A) is only partially refined, and an iterative refinement of each layer.
Still in the context of learning paradigms, it is important to address the case in which an initial model is built analytically and then refined through optimization algorithms. The most common solutions are optimizing parameters of the analytical formulation, or building a new adjustment layer that modifies the initial model’s output to achieve the target behavior. The difference between the original model and the desired outcome is often called residuals. This decomposition technique is common in robotic environments, particularly when analytical models struggle to accurately represent the complex array of physical interactions [14; 18; 22; 49; 60].
In a cooperative-competitive soccer environment, addressing the game’s multi-agent nature is crucial, even when developing low-level skills. This is particularly important when multiple agents contend for the ball and collisions are likely to occur. Multi-agent reinforcement learning (MARL) presents three major information structures [61]: a centralized setting where a central controller shares information with all agents, a decentralized setting where agents can communicate with each other, and a fully decentralized setting where no explicit communication occurs. The mathematical framework underlying a 3DSSL game can be characterized as a decentralized partially observable MDP (Dec-POMDP) [6]. However, during training, each agent can access all available information, making the widely-adopted centralized-learning-decentralized-execution scheme [15; 32; 50; 51] a desirable option.
@@ 3.2 Practical applications
Gao et al. [13] present an extensive list of skills and other relevant research concerning the 3DSSL. In the scope of this work, the main skill-related contributions can be classified into three categories — locomotion, dribbling, and kicking. For locomotion, most works focus on building a strong analytical foundation based on the Zero Moment Point (ZMP) and a model that abstracts the robot’s dynamics, both historically [16; 20] and in recent contributions [21; 30; 45; 46]. One popular simplification approach is the linear inverted pendulum model (LIPM).
The effectiveness of these models can be improved through various approaches, such as interpolation of keyframes [38; 47; 48; 56] or optimization methods. Optimization can be used to complement analytical models, either in parallel [23; 24; 35; 54] or sequentially [12]. It can be used over primitives extracted from observed trajectories [7] or soccer motion capture data [31], or to learn from scratch, as demonstrated in activities like walking [55] and running [2; 3; 36; 37].
Dribbling is an extension of locomotion, with the added effort of controlling the ball. This skill can be divided into close control dribble [2], involving precise ball manipulation while keeping it near the player’s feet; and speed dribble [28; 29; 40], where the robot executes precise kicks to send the ball a bit further and then swiftly pursues it.
Finally, in the literature, kicking is predominantly categorized into two main forms: as the culmination of locomotion skills (implying that the robot is moving at a non-negligible speed before kicking) [4; 43] or as a skill that involves a slower preparation stage before kicking. In the latter case, contributions vary from attempting to cover the largest distance on a given axis [9; 17; 19], including sideways kicks [10], to customizing the target distance [1; 26] and the kick direction [52]. The static or moving initial conditions represent a trade-off between accuracy and reaction time.
@@ 4 Skill-set-primitives
This section introduces a type of analytical structure that serves as the foundation for learning more complex behaviors. In this work, we propose skill-set-primitives, a subclass of motion primitives [8; 41] designed for a set of skills instead of a single behavior. This type of primitive captures the commonalities of a set of actions. For instance, while walking, humans lift each foot alternately, regardless of the direction of motion, even during rotations. This underlying movement template can also be applied to marching, running, or dribbling a ball.
@@
  • 1-to-many scheme — instead of using primitives as building blocks to develop other skills in a 1-to-1 or many-to-1 scheme [7; 31], we use them as an integrating foundation to develop and unite a specific set of skills;
@@ 5.2 Locomotion set
The skill-set-primitive was then frozen, and all parameters were reused for dribbling and pushing, except for $p_{y}$, which was increased to motivate a wider leg stance, except for robot R3 due to its default wide stance. See Appendix A.1, Table 2 for the optimized values. Note that $c_{z}$ is given as a percentage of the COM’s height when the robot is standing with straight legs, since different robot types vary in height and weight distribution. The second dribbling version reverts to the $p_{y}$ used during walking, as a wider leg stance is more likely to trigger ball-holding fouls under the new rules.
@@
Fig. 4 presents an execution sequence example, where the Step Baseline is active during walking, pushing, and dribbling. A matrix of skill transitions is provided in Appendix A.1, Table 3. Transitions were either trained as part of skill optimization, innate, if they emerged naturally, or assisted, when a gradual return to the Step Baseline is required before walking (as seen in Fig. 4, (a)).
@@ 6 Training Environment
This section describes the most important characteristics of the environments used to train the aforementioned skills. For a comprehensive list of environment characteristics, please refer to Appendix A.2 and A.3, which cover details such as state space, symmetry operations (referenced with index/multiplier notation [5]), action space modifications, episode setup, reward functions, and hyperparameters.
@@ 6.1 Sprint-Kick
The Sprint-Kick training environment is outlined in Fig. 5. A 2-hidden-layer deep neural network is trained with PPO and abstracted from the left and right sides of the robot. To do this, we relabel states and actions so that the learning policy is only aware of a front and a back leg. The robot initiates sprinting with the left leg in the front, and whenever the right leg takes the lead, all observations and actions are mirrored along the robot’s sagittal plane. The resulting action goes through a post-processing stage including a low-pass filter, and mapping raw joint values to their respective operational ranges. The primitive that captures the Sprint Forward motion cycle is then added to the subsequent Curved Sprint and Kick skills.
@@ 7 Results and Discussion
Demonstrations of all skills can be found on YouTube  [^5] https://www.youtube.com/playlist?list=PLIeNX3I5JIATFlOXfM93d6cijDdRheyyO For a comprehensive list of available videos, refer to Appendix A.5.. Major results are summarized in Table 1. The Sprint Forward skill, which is the base of the Sprint-Kick, attains an average speed of 3.69 m/s after stabilizing — an improvement over prior work [3] (2.5 m/s). While Melo et al. [36], reported an average speed of 3.76 m/s under controlled conditions, our sprint was the fastest sprint to be integrated into a team and demonstrated in an official RoboCup competition, in 2021. The Sprint-Kick sacrifices some speed for maneuverability, being ultimately used to score goals during kickoff set plays. It was also the first kick in the league to be performed while running. For more results, see Appendix A.4.
@@
Omnidirectional Walk maximum speed 0.70-0.90 m/s a
Kick Short Kick avg. error (3/6/9 m) 0.14/0.25/0.62 m
Long Kick average error 1.85 m
@@
Dribble v1 average fwd. speed 1.25-1.41 m/s b
@@
a depends on robot type and walking direction                   b depends on robot type
The Omnidirectional Walk was trained individually for each robot type, offering fast forward walking, rotating while walking in any direction, and precise approaches to kick the ball or wait in strategic positions. Maximum speeds range from 0.70 to 0.90 m/s, depending on walk direction and robot type. Comparing walking skills is challenging due to high variance in versatility. In 2015, Wang et al. [56] achieved 0.5 m/s for forward and backward walking, and 0.4 m/s for side walking, while retaining omnidirectional abilities. In the same year, Shafii et al. optimized a forward walk (0.78 m/s) and a side walk (0.48 m/s). It would be unfair to compare these speeds with an omnidirectional walk. Recent advancements, such as Kasaei et al.’s [23] forward walk with rotation (0.956 m/s) and Fischer et al.’s [12] specialization for robot type 4 (toes) with speeds of 1.3 m/s forward and 1.03 m/s backward, concern specific conditions. However, these approaches lack the all-encompassing control of an omnidirectional walk, with the ability to accelerate, decelerate, rotate, and change walking direction seamlessly and automatically in a single skill. Our behavior is best analyzed in action  [^6] Omnidirectional Walk training environment and demonstration: https://youtu.be/dXzIuZlOFZc.
The Short Kick offers variable distance but was tested at 3, 6, and 9 meters, with 1000 samples, to measure distance errors (see Table 1). The Long Kick originated a heatmap with a spatial distribution of final ball positions using robot type 1 (see Fig. 8). Reaching 20 m [9] or more [17] is not uncommon in the league. Moreover, some kicks have special features such as multi-directional targets [52], kick while moving [4; 43], or handling fast-moving balls, which magmaOffenburg mastered for the RoboCup 2023 Kick Rolling Ball Challenge. However, success depends on factors such as kick preparation quality and repeatability, duration, accuracy, and adaptability to in-game disruptions.
@@
In previous work [2], we introduced a dribble developed from scratch through RL, where the ball is carried between the robot’s legs at 0.31 m/s, which was never integrated into the team for being slow. Dribble v1 was the first close control dribble to be introduced in the 3D Soccer Simulation League, reaching 1.25 to 1.41 m/s, depending on the robot type. The second version complies with new rules by keeping the ball further than 12 cm, albeit at the cost of speed [^7] Dribble evolution, training environment, and demonstration: https://youtu.be/8UED_Zl-nbQ. Other dribble behaviors are not comparable [28; 29; 40], as they focus on kicking the ball further and rapidly pursuing it, rather than maintaining close control throughout.

### v2

@@ Abstract
The RoboCup 3D Soccer Simulation League serves as a competitive platform for showcasing innovation in autonomous humanoid robot agents through simulated soccer matches. Our team, FC Portugal, developed a new codebase from scratch in Python after RoboCup 2021. The team’s performance relies on a set of skills centered around novel unifying primitives and a custom, symmetry-extended version of the Proximal Policy Optimization algorithm. Our methods have been thoroughly tested in official RoboCup matches, where FC Portugal has won the last two main competitions, in 2022 and 2023. This paper presents our training framework, as well as a timeline of skills developed using our skill-set-primitives, which considerably improve the sample efficiency and stability of skills, and motivate seamless transitions. We start with a significantly fast sprint-kick developed in 2021 and progress to the most recent skill set, including a multi-purpose omnidirectional walk, a dribble with unprecedented ball control, a solid kick, and a push skill. The push addresses low-level collision scenarios and high-level strategies to increase ball possession. We address the resource-intensive nature of this task through an innovative multi-agent learning approach. Finally, we release the team’s codebase to the RoboCup community, providing other teams with a robust and modern foundation upon which they can build new features.
@@ 1 Introduction
RoboCup is an international competition standing at the forefront of robotics research and innovation. It features four main leagues: @Home for domestic robot assistance, Industrial for logistics and warehousing systems, Rescue for disaster strategies and rescue operations, and Soccer, the largest league, focusing on competitive robotics.
The primary objective of this project is to enable autonomous humanoid teams to win against human soccer champions by 2050 [31]. Achieving this goal requires addressing the disparities between human and robotic capabilities in decision-making, coordination, and skill execution.
Our focus lies on the 3D Soccer Simulation League (3DSSL) — a simulated soccer competition involving two teams of autonomous humanoid robots. Established in 2004, the league adopted the NAO robot model in 2008. Initially, matches were played on a 12 m by 8 m field with teams of three robots [9]. In 2012, the league expanded to its current dimensions of 30 m by 20  m, allowing for full teams of 11 players.
Some of the 3DSSL’s main challenges and research efforts have been directed toward skill optimization, real-time decision-making, and cooperative multi-agent strategies [16]. Robots must execute fundamental tasks like locomotion and kicking while dynamically adapting to game scenarios involving role allocation and team coordination. Additionally, addressing these issues also requires overcoming challenges such as sensor noise, localization under uncertainty, and collision-free path planning. Some of these concerns are also applicable to real-world robotics, giving the research in this area broader implications beyond robotic soccer or RoboCup.
This paper introduces a novel approach to develop a cohesive structure of skills that works in harmony at different levels of abstraction — from motor control to strategy — using primitives that represent the commonalities of a skill set. All the methods detailed in this paper have been validated through participation in several official competitions, serving as the central pillar of our team, FC Portugal.
Figure 1: Timeline of skills (based on skill-set-primitives) that were used in official competitions. Each rhombus indicates the tally of goals directly attributed to each skill in the respective competition
Figure 1 depicts a timeline of skills developed with skill-set-primitives. In 2020, we developed the Sprint-Kick  [^1] First Sprint-Kick demonstration in 2020 https://youtu.be/Yy1yCM5hwZI, a behavior that sprinted toward the ball during kickoff, resulting in 22% of our team’s goals in the Portuguese Robotics Open (PRO) 2021. In RoboCup (RC) 2021 (virtual event), the kickoff was refined, accounting for 38% of all our goals  [^2] Successful kickoffs in RoboCup 2021 https://youtu.be/3MND8RVUPBQ. We secured 3rd and 6th place, respectively, having good individual skills but lacking cohesion.
@@
The dominant dribbling skill allowed any player, even the goalkeeper, to quickly bypass all opponents and score, while keeping the ball between both feet, making defense nearly impossible without fouling. To improve gameplay, RoboCup 2023 (Bordeaux) introduced new rules to prevent ball-holding  [^3] Official rules are found at https://ssim.robocup.org/3d-simulation/3d-rules/, rendering our Dribble and Push skills unusable. Due to time constraints, only the Dribble was retrained to comply with the new rules. FC Portugal won the main competition for the third time in 2023.
@@ 2.2 Learning algorithm
The reinforcement learning algorithm used in this work is the Proximal Policy Optimization (PPO) [50] extended with Proximal Symmetry Loss (PSL) [^4] PPO+PSL algorithm: https://github.com/m-abr/Adaptive-Symmetry-Learning [6]. PPO was chosen due to its good performance in single-agent and cooperative multi-agent games [66]. PSL leverages the symmetry properties of the NAO robot (sagittal plane), to improve the training sample efficiency and also promote a more human-like behavior, without hindering asymmetric exploration. The extended objective function can be written as:
@@
$L^{PSL}(\theta,\omega)$ $=\hat{\mathbb{E}}_{t}\left[\,\text{{w}}_{\pi}\cdot L_{t}^{\pi}(\theta)+\text{{w}}_{V}\cdot L_{t}^{V}(\omega)\,\right],$ (6)
@@
where $\textit{w}_{\pi}$ and $\textit{w}_{V}$ are weight vectors, $V_{\omega}(f(s_{t}))$ is the symmetric state value, $V_{t}^{\text{targ}}$ is the same target value used by PPO in $L^{VF}$, and $x_{t}(\theta)$ is a symmetry probability ratio. An extensive description of this algorithm can be found in previous work [6].
@@
The RoboCup 3D simulation league uses SimSpark [64] as its physical multi-agent simulator. The league’s environment is a 30 m by 20 m soccer field containing several landmarks that can be used for self-localization: goalposts, corner flags, and lines. Each team consists of 11 humanoid robots modeled after the NAO robot, created by SoftBank Robotics. Agents get internal data (joints, accelerometer, gyroscope, and foot pressure sensors) with a one-step delay every 0.02 s, and visual data (restricted to a 120 vision cone) every 0.04 s (the visual data interval was 0.06 s prior to 2023). Agents can send joint commands every 0.02 s and message teammates every 0.04 s. In addition to the standard NAO model (type 0), there are alternative types with longer limbs (type 1), quicker feet (type 2), wider hips and longest limbs (type 3), and added toes (type 4).
@@ 3 Related Work
RoboCup is an international robotics competition that drives AI and robotics research, featuring standardized challenges for robotic teams. Most of the original challenges in 1997 [32] still stand today in the 3D soccer simulation league (3DSSL), including multi-agent collaboration, real-time reasoning, handling dynamic environments, and learning complex tasks.
@@ 3.1 Learning paradigms
Developing a team involves the creation of an interplay of skills at different levels of abstraction, requiring seamless transitions and coordination between objectives of varying granularity. Hierarchical reinforcement learning is one way of tackling this issue by allowing the decomposition of complex tasks. In light of this, Stone et al. [59] proposed the Layered Learning (LL) paradigm, where tasks are learned in a bottom-up approach, originally in a sequential way. A lower-level layer (A) is learned and frozen before learning the next layer (B). Posterior formulations include concurrent [63] and overlapping variations [39; 40]. In concurrent LL, (A) is initially learned in isolation, being later refined while simultaneously learning (B). In overlapping LL, there are 3 paradigms that can be succinctly described as: the combination of independently learned layers, a partial version of concurrent LL where (A) is only partially refined, and an iterative refinement of each layer.
Still in the context of learning paradigms, it is important to address the case in which an initial model is built analytically and then refined through optimization algorithms. The most common solutions are optimizing parameters of the analytical formulation, or building a new adjustment layer that modifies the initial model’s output to achieve the target behavior. The difference between the original model and the desired outcome is often called residuals. This decomposition technique is common in robotic environments, particularly when analytical models struggle to accurately represent the complex array of physical interactions. In this context, RL has been employed to refine control models, enhancing performance in manipulation and control tasks [22; 55]. Furthermore, applying residual adjustments has been shown to improve robotic locomotion, enabling more agile and stable movements in unpredictable terrains [26; 33; 65]. These techniques have been widely used, ranging from dexterous hand-object interactions [17] to improving grasping and throwing accuracy [67], highlighting the effectiveness of residual learning in refining robotic behaviors across diverse tasks.
In a cooperative-competitive soccer environment, addressing the game’s multi-agent nature is crucial, even when developing low-level skills. This is particularly important when multiple agents contend for the ball and collisions are likely to occur. Multi-agent reinforcement learning (MARL) presents three major information structures [68]: a centralized setting where a central controller shares information with all agents, a decentralized setting where agents can communicate with each other, and a fully decentralized setting where no explicit communication occurs. The mathematical framework underlying a 3DSSL game can be characterized as a decentralized partially observable MDP (Dec-POMDP) [8]. However, during training, each agent can access all available information, making the widely-adopted centralized-learning-decentralized-execution scheme [7; 19; 38; 56; 57] a desirable option.
@@ 3.2 Practical applications
Gao et al. [16] present an extensive list of skills and other relevant research concerning the 3DSSL. In the scope of this work, the main skill-related contributions can be classified into three categories — locomotion, dribbling, and kicking. For locomotion, most works focus on building a strong analytical foundation based on the Zero Moment Point (ZMP) and a model that abstracts the robot’s dynamics, both historically [20; 24] and in recent contributions [25; 36; 51; 52]. One popular simplification approach is the linear inverted pendulum model (LIPM).
The effectiveness of these models can be improved through various approaches, such as interpolation of keyframes [44; 53; 54; 62] or optimization methods. Optimization can be used to complement analytical models, either in parallel [27; 28; 41; 60] or sequentially [15]. It can be used over primitives extracted from observed trajectories [10] or soccer motion capture data [37], or to learn from scratch, as demonstrated in activities like walking [61] and running [2; 3; 42; 43].
Dribbling is an extension of locomotion, with the added effort of controlling the ball. This skill can be divided into close control dribble [2], involving precise ball manipulation while keeping it near the player’s feet; and speed dribble [34; 35; 46], where the robot executes precise kicks to send the ball a bit further and then swiftly pursues it.
Finally, in the literature, kicking is predominantly categorized into two main forms: as the culmination of locomotion skills (implying that the robot is moving at a non-negligible speed before kicking) [5; 49] or as a skill that involves a slower preparation stage before kicking. In the latter case, contributions vary from attempting to cover the largest distance on a given axis [12; 21; 23], including sideways kicks [13], to customizing the target distance [1; 30] and the kick direction [58]. The static or moving initial conditions represent a trade-off between accuracy and reaction time.
@@ 4 Skill-set-primitives
This section introduces a type of analytical structure that serves as the foundation for learning more complex behaviors. In this work, we propose skill-set-primitives, a subclass of motion primitives [11; 47] designed for a set of skills instead of a single behavior. This type of primitive captures the commonalities of a set of actions. For instance, while walking, humans lift each foot alternately, regardless of the direction of motion, even during rotations. This underlying movement template can also be applied to marching, running, or dribbling a ball.
@@
  • 1-to-many scheme — instead of using primitives as building blocks to develop other skills in a 1-to-1 or many-to-1 scheme [10; 37], we use them as an integrating foundation to develop and unite a specific set of skills;
@@ 5.2 Locomotion set
The skill-set-primitive was then frozen, and all parameters were reused for dribbling and pushing, except for $p_{y}$, which was increased to motivate a wider leg stance, except for robot R3 due to its default wide stance. Refer to A.1, Table 8 for the optimized values. Note that $c_{z}$ is given as a percentage of the COM’s height when the robot is standing with straight legs, since different robot types vary in height and weight distribution. The second dribbling version reverts to the $p_{y}$ used during walking, as a wider leg stance is more likely to trigger ball-holding fouls under the new rules.
@@
Fig. 4 presents an execution sequence example, where the Step Baseline is active during walking, pushing, and dribbling. A matrix of skill transitions is provided in A.1, Table 9. Transitions were either trained as part of skill optimization, innate, if they emerged naturally, or assisted, when a gradual return to the Step Baseline is required before walking (as seen in Fig. 4, (a)).
@@ 6 Training Environment
This section describes the most important characteristics of the environments used to train the aforementioned skills. For a comprehensive list of environment characteristics, please refer to A.2 and A.3, which cover details such as state space, symmetry operations (referenced with index/multiplier notation [6]), action space modifications, episode setup, reward functions, and hyperparameters.
@@ 6.1 Sprint-Kick
The Sprint-Kick training environment is outlined in Fig. 5. A 2-hidden-layer deep neural network is trained with PPO and abstracted from the left and right sides of the robot. To do this, we relabel states and actions so that the learning policy is only aware of a front and a back leg. The robot begins sprinting with the left leg in the front, and whenever the right leg takes the lead, all observations and actions are mirrored along the robot’s sagittal plane. The resulting action goes through a post-processing stage including a low-pass filter, and mapping raw joint values to their respective operational ranges. The primitive that captures the Sprint Forward motion cycle is then added to the subsequent Curved Sprint and Kick skills.
@@ 7 Results and Discussion
Demonstrations of all skills can be found on YouTube  [^5] https://www.youtube.com/playlist?list=PLIeNX3I5JIATFlOXfM93d6cijDdRheyyO For a comprehensive list of available videos, refer to Appendix A.5.. Major results are summarized in Table 1. The Sprint Forward skill, which is the base of the Sprint-Kick, attains an average speed of 3.69 m/s after stabilizing — an improvement over prior work [3] (2.5 m/s). While Melo et al. [42], reported an average speed of 3.76 m/s under controlled conditions, our sprint was the fastest sprint to be integrated into a team and demonstrated in an official RoboCup competition. The Sprint-Kick sacrifices some speed for maneuverability, being ultimately used to score goals during kickoff set plays. It was also the first kick in the league to be performed while running. For more results, see Appendix A.4.
@@
Omnidirectional Walk maximum speed 0.70--0.90 m/s 111depends on robot type and walking direction
Short Kick avg. error (3/6/9 m) 0.14/0.25/0.62 m
Kick Long Kick average error 1.85 m
@@
Dribble v1 average fwd. speed 1.25--1.41 m/s 222depends on robot type
@@
The Omnidirectional Walk was trained individually for each robot type, offering fast forward walking, rotating while walking in any direction, and precise approaches to kick the ball or wait in strategic positions. Maximum speeds range from 0.70 to 0.90 m/s, depending on walk direction and robot type. Comparing walking skills is challenging due to high variance in versatility. In 2015, Wang et al. [62] achieved 0.5 m/s for forward and backward walking, and 0.4 m/s for side walking, while retaining omnidirectional abilities. In the same year, Shafii et al. [52] optimized a forward walk (0.78 m/s) and a side walk (0.48 m/s). It would be unfair to compare these speeds with an omnidirectional walk. Recent advancements, such as Kasaei et al.’s [27] forward walk with rotation (0.956 m/s) and Fischer et al.’s [15] specialization for robot type 4 (toes) with speeds of 1.3 m/s forward and 1.03 m/s backward, concern specific conditions. However, these approaches lack the all-encompassing control of an omnidirectional walk, with the ability to accelerate, decelerate, rotate, and change walking direction seamlessly and automatically in a single skill. Our behavior is best analyzed in action  [^6] Omnidirectional Walk training environment and demonstration: https://youtu.be/dXzIuZlOFZc.
The Short Kick offers variable distance but was tested at 3, 6, and 9 meters, with 1000 samples, to measure distance errors (see Table 1). The Long Kick originated a heatmap with a spatial distribution of final ball positions using robot type 1 (see Fig. 8). Reaching 20 m [12] or more [21] is not uncommon in the league. Moreover, some kicks have special features such as multi-directional targets [58], kick while moving [5; 49], or handling fast-moving balls, which magmaOffenburg mastered for the RoboCup 2023 Kick Rolling Ball Challenge. However, success depends on factors such as kick preparation quality and repeatability, duration, accuracy, and adaptability to in-game disruptions.
@@
In previous work [2], we introduced a dribble developed from scratch through RL, where the ball is carried between the robot’s legs at 0.31 m/s, which was never integrated into the team for being slow. Dribble v1 was the first close control dribble to be introduced in the 3D Soccer Simulation League, reaching 1.25 to 1.41 m/s, depending on the robot type. The second version complies with new rules by keeping the ball further than 12 cm, albeit at the cost of speed [^7] Dribble evolution, training environment, and demonstration: https://youtu.be/8UED_Zl-nbQ. Other dribble behaviors are not comparable [34; 35; 46], as they focus on kicking the ball further and rapidly pursuing it, rather than maintaining close control throughout.